Calculate the following:
Quantity
Total Fixed
Cost
Total Variable
Cost
Total
Cost
Average
Cost
Marginal

25 10 18 ? ? ?
26 10 20 ? ? ?
27 10 21 ? ? ?
Analyse the changes in the calculated costs as quantity produced increases.
1
Expert's answer
2017-12-05T12:47:06-0500
TC = FC + VC,
ATC = TC/Q,
MC = (TC2 - TC1)/(Q2 - Q1).
Q FC VC TC ATC MC
25 10 18 28 1.12 -
26 10 20 30 1.154 2
27 10 21 31 1.148 1